OpenStudy (anonymous):

A right triangle has legs with lengths of 5 cm and 12 cm. Find the length of the hypotenuse.

OpenStudy (31356):

Go to this? http://www.wikihow.com/Find-the-Length-of-the-Hypotenuse

OpenStudy (31356):

A^2+B^2=C^2

OpenStudy (31356):

As a result I got 13?

OpenStudy (31356):

5^2+12^2=25+144=169 find square root and you get 13!!